C = 160 + 0.6 * YD

I = 150

G = 150

T = 100

Solve for the following variables. a) Equilibrium GDP, Y. b) Disposable income, YD. vs)

Consumer expenditure, C.


1
Expert's answer
2021-02-10T07:09:14-0500

At equilibrium;

Y=AD

C = 160 + 0.6 * YD

I = 150

G = 150

T = 100

Thus;

Ye=160 + 0.6 * YD+150+15

But YD=disposable income

               =Income-tax

               =150-100

YD=50

Ye=160+0.6(50) +150+15

               =160+3+150+15

=328

Therefore, Consumer expenditure

C=160+0.6*YD

=160+0.6(50)

               =168